Which of the following is not a Boolean identity
A + 0 =0
A X A = A
A + A = A
A + 1 = 1
1
Expert's answer
2016-02-16T00:00:57-0500
The one of Boolean Additive Identities is
A+0=A.
So, A + 0 =0 is not a Boolean Identity.
Answer: A + 0 =0.